Write and solve two mathmatical sentences for the following problem:
The length of a crayon box base is 2 more than 3 times its width. If the perimeter of the base of the crayon box is 14 centimeters, find the length and width.

Answer by rapaljer(4671)

Let x= width
3x+2 = length

The formula for Perimeter is
2W + 2L = P
2(x)+2(3x+2)=14
2x+6x+4=14
8x+4=14
8x=10
x=10/8=5/4 cm
Length = 3x+2
Length = 3(5/4)+2
Length = 15/4 +2 =15/4 + 8/4=23/4 cm

Check: 2W + 2L
2(5/4) + 2(23/4)=10/4 + 46/4= 56/4= 14 cm. It checks!!

Please see my own website for additional explanation on perimeter problems. Do a Google search for my name "dr robert j rapalje," and look for "Rapalje Homepage" near the top of the list. Then find the link "Basic, Intermediate, and College Algebra: One Step at a Time." Select "Basic Algebra," and look in Chapter 1 for Section 1.11. You will find a complete explanation with examples and exercises, with many of the exercises solved in the MATH IN LIVING COLOR pages that are attached.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus